# Unsupervised Model Learning for Quantum Impurity Systems

### v1

@@ Abstract
Generalized quantum impurity models – which feature a few localized and strongly-correlated degrees of freedom coupled to itinerant conduction electrons – describe diverse physical systems, from magnetic moments in metals to nanoelectronics quantum devices such as quantum dots or single-molecule transistors. Correlated materials can also be understood as self-consistent impurity models through dynamical mean field theory. Accurate simulation of such models is challenging, especially at low temperatures, due to many-body effects from electronic interactions, resulting in strong renormalization. In particular, the interplay between local impurity complexity and Kondo physics is highly nontrivial. A common approach, which we further develop in this work, is to consider instead a simpler effective impurity model that still captures the low-energy physics of interest. The mapping from bare to effective model is typically done perturbatively, but even this can be difficult for complex systems, and the resulting effective model parameters can anyway be quite inaccurate. Here we develop a non-perturbative, unsupervised machine learning approach to systematically obtain low-energy effective impurity-type models, based on the renormalization group framework. The method is shown to be general and flexible, as well as accurate and systematically improvable. We benchmark the method against exact results for the Anderson impurity model, and provide an outlook for more complex models beyond reach of existing methods.
@@
As an example, it is often more convenient to analyse an effective lattice model in second quantized form than the ab initio treatment involving an all-orbital description of the constituent atoms [2, 3]. Methods such as the constrained random-phase approximation [4, 5], coupled cluster downfolding [6], density matrix downfolding [7] or the Pariser-Parr-Pople model for molecules [8, *pariser1953semi2] were devised to systematically eliminate inactive degrees of freedom and account for them with renormalized parameters of a reduced model.
@@
Figure 1: Schematic of the simplest bare (blue) and effective (orange) impurity models, together with an illustrative comparison of their impurity entropies after optimization.
@@
These methods have in common that they require some prior knowledge about the bare model to derive the effective model. This situation shares some similarity with that of the inverse problem [10] encountered in statistical inference or machine learning [11], that seeks to infer a probabilistic model from observed data. Recently machine learning techniques have been explored to construct effective Hamiltonians from simulated data [12, 13, 14, 15] or from experimentally measured data [16, 17, 18]. Another field where the inference of Hamiltonians has gained considerable relevance, is in the analogue simulation of quantum Hamiltonians on quantum hardware [19, *georgescu2014quantum, 21, *hensgens2017quantum, 23]. Indeed, machine learning inspired process- and Hamiltonian-tomography methods have been developed to infer the precise Hamiltonian that NISQ devices actually simulate, given hardware imperfections [24, 25, 26, 27, 28, 29].
@@
To derive a good effective Hamiltonian, it is not enough to infer a model with the desired properties – it must also be meaningfully simpler than the bare model. The canonical method to systematically eliminate degrees of freedom and obtain effective models is the renormalization group (RG) [30, 31, 32]. On the level of the thermal density matrix a single RG transformation acts as the partial trace over the high energy degrees of freedom [33]. In the context of quantum physics, a central concept from information theory is that thermal states encode the corresponding Hamiltonian [34], such that comparing the ensemble of thermal states from two systems is equivalent to comparing the two Hamiltonians. This provides a way to measure the ‘difference’ between a Hamiltonian before and after renormalization. By minimizing the distance in Hamiltonian space, one can therefore in principle optimize a simplified effective model to best approximate some more complicated bare Hamiltonian after renormalization. We rigorously derive such an approach in this paper for a specific class of systems known as ‘quantum impurity’ models [35]. These models involve localized interacting quantum degrees of freedom coupled to one or more continuum baths of noninteracting conduction electrons. Generalized quantum impurity models comprise the microscopic description of semiconductor quantum dot devices [36] and complex single-molecule junctions [37]. They also underpin the theoretical understanding of correlated materials through dynamical mean field theory [38]. Developing a strategy for systematically and accurately deriving simple effective models to better understand and simulate complex systems is the ultimate application of this work.
@@
In particular, a prerequisite for the design and use of complex quantum nanoelectronics devices with advanced functionality beyond the classical paradigm, is a fundamental understanding of their low-temperature correlated electron physics and quantum transport properties. However, this is a notoriously difficult theoretical challenge because of the subtle interplay between the orbital and spin complexity of the nanostructure, determined by its structure and chemistry; strong electron interactions due to quantum confinement; and the coupling to $\sim 10^{23}$ conduction electrons in the external circuit. This results in nontrivial quantum phenomena such as Coulomb blockade [39], various forms of Kondo effect [39, 40, 37], and quantum interference [41, 42] – all of which strongly affect low-temperature electronic conductance through the device, and hence its functionality. As with coupled quantum dot devices [43, 44, *mitchell2016universality, 46], entangled spin and charge degrees of freedom can give rise to new physics in single molecule junctions. It is therefore a formidable task to derive simplified effective models that can still describe this range of physics in these kinds of quantum device.
@@
A perturbative approach to this problem in the context of molecular electronics [37] maps the microscopic model of a single-molecule junction to an effective two-channel Kondo model. This method captures simultaneously the effect on quantum transport from quantum interference and Kondo physics – but is inevitably approximate. Since thermodynamic observables flow under RG and the bare and effective model are defined at different energy scales, it is also not a priori clear whether such minimal Kondo models are sufficiently general to reproduce local observables of interest in the bare model [47]. In this paper we show how an RG analysis of effective interaction terms can be used to determine thermodynamic observables that are comparable across different energy scales. This allows us to introduce a novel machine learning (ML) methodology to derive accurate effective models for complex quantum impurity problems, that works by optimizing generalized (minimally constrained) models, and ensuring that local observables are correctly reproduced. We show by way of explicit examples that the low-energy Kondo physics is simultaneously captured. The parameters of the effective model are optimized by minimizing the Kullback-Leibler divergence (KLD) [48] that compares its ensemble of thermal states to that of the bare model. Information on the target is extracted from a numerical simulation of the system of interest. However, the full solution of the bare model is not required: minimization of the KLD requires only an estimation of thermal expectation values corresponding to specific local operators at relatively high temperatures. This can be achieved with any suitable quantum impurity solver [49, 50, 51, 52, *stadler2016interleaved]. Furthermore, the KLD can be shown to be convex under reasonable assumptions (see Appendix A) and its gradient is known analytically in closed form [54], which makes it an ideal optimization problem. We refer to this method as unsupervised model learning (UML). We demonstrate the efficacy of the UML method by application to the Anderson impurity model (AIM) [35] and obtain some new non-perturbative results for this old problem. Finally we give an outlook to the application of this framework to more complex problems, where the development of tractable effective models is essential for the study of nontrivial correlated electron physics at low temperatures.
I Method
@@
The generalized KLD quantifies how distinguishable $\hat{\rho}_{1}$ is from $\hat{\rho}_{2}$. The thermal density matrix $\hat{\rho}=\frac{1}{\mathcal{Z}}e^{-\beta\hat{H}}$ is fully defined by the system Hamiltonian $\hat{H}$ (and inverse temperature $\beta$). Thus, the thermal density matrix can be seen as a proxy for its defining Hamiltonian, and the KLD as a measure of distinguishability between two Hamiltonians. To emphasize this we denote the KLD for two thermal density matrices $\hat{\rho}_{1}=\frac{1}{\mathcal{Z}_{1}}e^{-\beta\hat{H}_{1}}$ and $\hat{\rho_{2}}=\frac{1}{\mathcal{Z}_{2}}e^{-\beta\hat{H}_{2}}$ as $D_{KL}[\hat{H}_{2}:\hat{H}_{1}]$.
@@
by minimizing $D_{KL}[\hat{H}_{\rm bare}:\hat{H}_{\rm eff}(\boldsymbol{\theta})]$ with respect to $\boldsymbol{\theta}$ for a set of operators $\{\hat{h}_{i}\}$. The minimization yields the optimal couplings $\boldsymbol{\theta}^{*}$ to represent $\hat{H}_{\rm{bare}}$ with $\hat{H}_{\rm{eff}}(\boldsymbol{\theta}^{*})$. To achieve this we represent the impurity thermal density matrix $\hat{\rho}_{\rm{imp}}=\operatorname{tr}_{\rm{bath}}\left[e^{-\beta\hat{H}}/\mathcal{Z}\right]$ as a classical probability distribution which can be used to evaluate the KLD. Using the hybridization perturbation expansion one can write the partition function of $\hat{H}_{\rm bare}$ as a summation of the weights of all impurity occupation diagrams $x$, which is subdivided into a summation over all corresponding diagrams in terms of impurity eigenstates $\{\alpha\}_{x}$ [57]
$\begin{gathered}\mathcal{Z}=\mathcal{Z}_{\rm bath}\sum_{x}\sum_{\{\alpha\}_{x}}w(\{\alpha\}_{x})\;,\\ w(\{\alpha\}_{x})=e^{-\langle\hat{H}^{\rm{imp}}\rangle_{\{\alpha\}_{x}}}\Lambda_{\{\alpha\}_{x}}\det(\Delta_{x})\;,\end{gathered}$ (5)
where $w(\{\alpha\}_{x})$ is the weight of a distinct Feynman diagram labelled by the eigenstate diagram $\{\alpha\}_{x}$, $\Delta_{x}$ is the antiperiodic hybridization matrix, $\Lambda_{\{\alpha\}_{x}}$ the sequence of impurity operators comprising the diagram $\{\alpha\}_{x}$ while being projected onto the eigenbasis of $\hat{H}^{\rm{imp}}$ and $\langle\hat{H}^{\rm{imp}}\rangle_{\{\alpha\}_{x}}$ is the average value of the impurity Hamiltonian over the diagram $\{\alpha\}_{x}$.
@@
$P(\{\alpha\}_{x})=(\mathcal{Z}_{\rm{bath}}/\mathcal{Z})w(\{\alpha\}_{x})\;,$
@@
where we have used the term admissible ($\rm ad$) to denote diagrams that involve eigenstates of $\hat{H}^{\rm imp}_{\rm bare}$ for which $\hat{H}^{\rm imp}_{\rm eff}$ has analogues eigenstates. To minimize the distinguishability between $\hat{H}_{\text{bare}}$ and $\hat{H}_{\rm{eff}}(\boldsymbol{\theta})$ we use gradient descent (GD) methods [11]. For this, we have the analytic gradient of the KLD in closed form
@@
where the admissibility operator $\hat{\Omega}_{\rm ad}$ connects the effective and bare Fock space, by mapping the effective eigenstates to analogous bare eigenstates (see Sec. II for a detailed derivation) and thus, eliminates all non-admissible states from the bare thermal average, while on the effective Hilbert space it holds that
$\langle\hat{\Omega}_{\rm ad}\rangle_{\rm{eff}}=\langle\hat{\mathbbm{1}}\rangle_{\rm{eff}}=1\;.$ (10)
From this form of the gradient (Eq. 9) it follows immediately that the minimum is found when the impurity observables match
@@
for all effective impurity operators $\hat{h}_{i}$ of the effective model (from now on we simplify the notation $\langle\hat{\Omega}_{\rm ad}^{\dagger}\hat{h}_{i}\hat{\Omega}_{\rm ad}\rangle_{\rm{bare}}\rightarrow\langle\hat{h}_{i}\rangle_{\rm bare}$). Thus follows that according to Eq. 8 the optimal low-energy effective model $\hat{H}_{\operatorname{\textrm{eff}}}$, for a given bare model $\hat{H}_{\rm bare}$, matches the thermodynamic expectation values of all effective impurity operators $\hat{h}_{i}$ in the bare model.
@@
The second order derivative of the KLD corresponds to the second order derivative of the effective free energy $\mathcal{F}_{\rm eff}(\boldsymbol{\theta})=-1/\beta\log\operatorname{tr}\left[\exp(-\beta\hat{H}_{\rm eff}(\boldsymbol{\theta}))\right]$ which, provided that all operators $\{\hat{h}_{i}\}$ mutually commute, becomes convex (as shown in Appendix A). It is important to note that the gradient obtained from the KLD Eq. 9 is equivalent to the gradient of the thermodynamic entropy, which is the loss-function in terms of the maximum entropy principle. The application of the maximum entropy principle to Hamiltonian tomography was recently proposed by Anshu et al. [24] in similar spirit as the here presented method with a great focus on convex problems, while here the focus lies on the RG analysis of the terms constituting $\hat{H}_{\rm eff}$.
@@
Given the convexity of Eq. 8 we can make some notes about the complexity of the optimization of Eq. 8. The complexity of the optimization depends on the method used to compute the observables comprising the gradient Eq. 9. The most expensive step is to calculate the observables $\langle\hat{h}_{i}\rangle_{\rm bare}$ in the bare model. However, this needs to be done just once (e.g. with QMC methods [58, 49]). On the other hand, multiple calculations of $\langle\hat{h}_{i}\rangle_{\rm eff}$ are required during the optimization. But these are performed on the simplified effective model, and are therefore inexpensive (they can be done with e.g. the NRG [50, 51, 52, *stadler2016interleaved] (NRG) method). For the results presented in the following, we have used CT-QMC and NRG to calculate the observables and the Adagrad or Adam GD method [59].
II Effective Hamiltonians

### v2

# Unsupervised Learning of Effective Quantum Impurity Models

@@ Abstract
Generalized quantum impurity models – which feature a few localized and strongly-correlated degrees of freedom coupled to itinerant conduction electrons – describe diverse physical systems, from magnetic moments in metals to nanoelectronics quantum devices such as quantum dots or single-molecule transistors. Correlated materials can also be understood as self-consistent impurity models through dynamical mean field theory. Accurate simulation of such models is challenging, especially at low temperatures, due to many-body effects from electronic interactions, resulting in strong renormalization. In particular, the interplay between local impurity complexity and Kondo physics is highly nontrivial. A common approach, which we further develop in this work, is to consider instead a simpler effective impurity model that still captures the low-energy physics of interest. The mapping from bare to effective model is typically done perturbatively, but even this can be difficult for complex systems, and the resulting effective model parameters can anyway be quite inaccurate. Here we develop a non-perturbative, unsupervised machine learning approach to systematically obtain low-energy effective impurity-type models, based on the renormalization group framework. The method is shown to be general and flexible, as well as accurate and systematically improvable. We benchmark the method against exact results for the Anderson impurity model, and provide an outlook for more complex models beyond the reach of existing methods.
@@
As an example, it is often more convenient to analyse an effective lattice model in second quantized form than the ab initio treatment involving an all-orbital description of the constituent atoms [2, 3, 4]. Methods such as the constrained random-phase approximation [5, 6], coupled cluster downfolding [7], density matrix downfolding [8] or the Pariser-Parr-Pople model for molecules [9, *pariser1953semi2] were devised to systematically eliminate inactive degrees of freedom and account for them with renormalized parameters of a reduced model.
@@
Figure 1: Schematic of the simplest bare (blue) and effective (orange) impurity models, together with an illustrative comparison of their impurity entropies after model optimization.
@@
These methods have in common that they require some prior knowledge about the bare model to derive the effective model. This situation shares some similarity with that of the inverse problem [11] encountered in statistical inference or machine learning [12], that seeks to infer a probabilistic model from observed data. Recently machine learning techniques have been explored to construct effective Hamiltonians from simulated data [13, 14, 15, 16] or from experimentally measured data [17, 18, 19]. Another field where the inference of Hamiltonians has gained considerable relevance, is in the analogue simulation of quantum Hamiltonians on quantum hardware [20, *georgescu2014quantum, 22, *hensgens2017quantum, 24]. Indeed, machine learning inspired process- and Hamiltonian-tomography methods have been developed to infer the precise Hamiltonian that NISQ devices actually simulate, given hardware imperfections [25, 26, 27, 28, 29, 30].
@@
A good effective Hamiltonian must not just reproduce desired properties of the bare model, it must also be meaningfully simpler than the bare model. The canonical method to systematically eliminate degrees of freedom and obtain effective models is the renormalization group (RG) [31, 32, 33]. On the level of the thermal density matrix a single RG transformation acts as the partial trace over the high energy degrees of freedom [34]. In the context of quantum physics, a central concept from information theory is that thermal states encode the corresponding Hamiltonian [35], such that comparing the ensemble of thermal states from two systems is equivalent to comparing the two Hamiltonians. This provides a way to measure the ‘difference’ between a Hamiltonian before and after renormalization. By minimizing the distance in Hamiltonian space, one can therefore in principle optimize a simplified effective model to best approximate some more complicated bare Hamiltonian after renormalization. We rigorously derive such an approach in this paper for a specific class of systems known as ‘quantum impurity’ models [36]. These models involve localized interacting quantum degrees of freedom (the ‘impurity’ $\hat{H}^{\rm imp}$) coupled (throguh $\hat{H}^{\rm hyb}$) to one or more continuum baths of noninteracting conduction electrons ($\hat{H}^{\rm bath}$), as described generically by:
$\hat{H}=\hat{H}^{\rm imp}+\hat{H}^{\rm bath}+\hat{H}^{\rm hyb}\;.$ (1)
Generalized quantum impurity models describe semiconductor quantum dot devices [37] and complex single-molecule junctions [38]. They also underpin our understanding of correlated materials through dynamical mean field theory [39]. Developing a strategy for systematically and accurately deriving simple effective models to better understand and simulate complex systems is the ultimate application of this work.
@@
In particular, a prerequisite for the design and use of complex quantum nanoelectronics devices with advanced functionality beyond the classical paradigm, is a fundamental understanding of their low-temperature correlated electron physics and quantum transport properties. However, this is a notoriously difficult theoretical challenge because of the subtle interplay between the orbital and spin complexity of the nanostructure, determined by its structure and chemistry; strong electron interactions due to quantum confinement; and the coupling to $\sim 10^{23}$ conduction electrons in the external circuit. This results in nontrivial quantum phenomena such as Coulomb blockade [40], various forms of Kondo effect [40, 41, 38], and quantum interference [42, 43] – all of which strongly affect low-temperature electronic conductance through the device, and hence its functionality. As with coupled quantum dot devices [44, 45, *mitchell2016universality, 47], entangled spin and charge degrees of freedom can give rise to new physics in single molecule junctions. It is therefore a formidable task to derive simplified effective models that can still describe this range of physics in these kinds of quantum device.
A basic illustrative example is given in Fig. 1. Here we show the temperature-dependence of a physical observable, in this case the entropy $S(T)$, for a bare quantum impurity model (blue), compared with that of a simpler effective model (orange). One can immediately identify that the effective model has lower complexity than the bare one, because in the high-temperature limit it has a lower entropy – signifying that the effective model has a smaller number of active degrees of freedom than the bare model. At high temperatures, physical properties of the two models are distinctly different. But below some low-energy scale, the behavior of the two models becomes essentially indistinguishable. The low-energy physics is perfectly captured by the simpler effective model. We will return to the specific details of this example later.
@@
In the context of molecular electronics, the complex microscopic models describing single-molecule junctions can be mapped to much simpler effective two-channel Kondo models using a perturbative approach, as described in Refs. [38, 43]. This method captures simultaneously the effect of quantum interference and Kondo physics – but is inevitably approximate due to the perturbative treatment. Since thermodynamic observables flow under RG and the bare and effective model are defined at different energy scales, it is also not a priori clear whether such minimal Kondo models are sufficiently general to reproduce local observables of interest in the bare model [48]. In this paper we show how an RG analysis of effective interaction terms can be used to determine thermodynamic observables that are comparable across different energy scales. This allows us to introduce a novel machine learning (ML) methodology to derive accurate effective models for complex quantum impurity problems, that works by optimizing generalized (minimally constrained) models, and ensuring that local observables are correctly reproduced. We show by way of explicit examples that the low-energy Kondo physics is simultaneously captured. The parameters of the effective model are optimized by minimizing the Kullback-Leibler divergence (KLD) [49] that compares its ensemble of thermal states to that of the bare model. Information on the target is extracted from a numerical simulation of the system of interest. However, the full solution of the bare model is not required: minimization of the KLD requires only an estimation of thermal expectation values corresponding to specific local operators at relatively high temperatures. This can be achieved with any suitable quantum impurity solver [50, 51, 52, 53, *stadler2016interleaved]. Furthermore, the KLD can be shown to be convex under reasonable assumptions (see Appendix A) and its gradient is known analytically in closed form [55], which makes it an ideal optimization problem. We refer to this method as ‘unsupervised model learning’ (UML). We demonstrate the efficacy of the UML method by application to the Anderson impurity model (AIM) [36] and obtain some new non-perturbative results for this old problem. Finally we give an outlook to the application of this framework to more complex problems, where the development of tractable effective models is essential for the study of nontrivial correlated electron physics at low temperatures.
I Overview of the Method
@@
The generalized KLD quantifies how distinguishable $\hat{\rho}_{1}$ is from $\hat{\rho}_{2}$. The thermal density matrix $\hat{\rho}=\frac{1}{\mathcal{Z}}e^{-\beta\hat{H}}$ is fully defined by the system Hamiltonian $\hat{H}$ (and inverse temperature $\beta$). Here $\mathcal{Z}={\rm tr}[e^{-\beta\hat{H}}]$ is the partition function of $\hat{H}$. Thus, the thermal density matrix can be seen as a proxy for its defining Hamiltonian, and the KLD as a measure of distinguishability between two Hamiltonians. To emphasize this we denote the KLD for two thermal density matrices $\hat{\rho}_{1}=\frac{1}{\mathcal{Z}_{1}}e^{-\beta\hat{H}_{1}}$ and $\hat{\rho_{2}}=\frac{1}{\mathcal{Z}_{2}}e^{-\beta\hat{H}_{2}}$ as $D_{KL}[\hat{H}_{2}:\hat{H}_{1}]$.
@@
by minimizing $D_{KL}[\hat{H}_{\rm bare}:\hat{H}_{\rm eff}(\boldsymbol{\theta})]$ with respect to the set of parameters $\{\theta_{i}\}$ corresponding to the operators $\{\hat{h}_{i}\}$. The minimization yields the vector of optimal couplings $\boldsymbol{\theta}^{*}$ to represent $\hat{H}_{\rm{bare}}$ with $\hat{H}_{\rm{eff}}(\boldsymbol{\theta}^{*})$. In general this is an extremely challenging problem.
Specializing now to quantum impurity models of the type Eq. 1, the KLD can be evaluated by representing the reduced thermal density matrix of the impurity after tracing out the bath, $\hat{\rho}_{\rm{imp}}=\operatorname{tr}_{\>\rm{bath}}\left[e^{-\beta\hat{H}}/\mathcal{Z}\right]$, as a classical probability distribution. The fact that $\hat{\rho}_{\rm{imp}}$ can be viewed as a classical probability distribution can be seen as follows. Using the hybridization perturbation expansion [58], one can write the partition function $\mathcal{Z}_{\rm bare}$ of $\hat{H}_{\rm bare}$ as a sum over the weights $w_{x}$ of all impurity occupation diagrams, denoted $x$. We further subdivide each such diagram $x$ into a sum over all corresponding diagrams $w(\{\alpha\}_{x})$ labelled in terms of impurity eigenstates $\{\alpha\}_{x}$,
$\begin{gathered}\mathcal{Z}=\mathcal{Z}^{\rm bath}\sum_{x}\sum_{\{\alpha\}_{x}}w(\{\alpha\}_{x})\;,\\ w(\{\alpha\}_{x})=e^{-\langle\hat{H}^{\rm{imp}}\rangle_{\{\alpha\}_{x}}}\Lambda_{\{\alpha\}_{x}}\det(\Delta_{x})\;,\end{gathered}$ (6)
where $w(\{\alpha\}_{x})$ is the weight of a distinct Feynman diagram labelled by the eigenstate diagram $\{\alpha\}_{x}$, $\Delta_{x}$ is the antiperiodic hybridization matrix [58], $\Lambda_{\{\alpha\}_{x}}$ the sequence of impurity operators comprising the diagram $\{\alpha\}_{x}$ while being projected onto the eigenbasis of $\hat{H}^{\rm{imp}}$ and $\langle\hat{H}^{\rm{imp}}\rangle_{\{\alpha\}_{x}}$ is the average value of the impurity Hamiltonian over the diagram $\{\alpha\}_{x}$. We have also defined $\mathcal{Z}^{\rm bath}$ as the partition function of the free (decoupled) bath $\hat{H}^{\rm bath}$. The details of the derivation and precise definitions of the terms appearing in Eq. 6 are given in Appendix A.
@@
$P(\{\alpha\}_{x})=(\mathcal{Z}^{\rm{bath}}/\mathcal{Z})w(\{\alpha\}_{x})\;,$
@@
where we have used the term admissible ($\rm ad$) to denote diagrams that involve eigenstates of $\hat{H}^{\rm imp}_{\rm bare}$ for which $\hat{H}^{\rm imp}_{\rm eff}$ has analogues eigenstates. This is important because the dimensionality of the impurity subspace is lower for the effective model than in the bare model, by construction. To minimize the distinguishability between $\hat{H}_{\text{bare}}$ and $\hat{H}_{\rm{eff}}(\boldsymbol{\theta})$ we use gradient descent (GD) methods [12]. For this, we need the analytic gradient of the KLD,
@@
where the admissibility operator $\hat{\Omega}_{\rm ad}$ connects the effective and bare Fock space, by mapping the effective eigenstates to analogous bare eigenstates (see Sec. II.1 for a detailed derivation and discussion). Thus, $\hat{\Omega}_{\rm ad}$ eliminates all non-admissible states from the bare thermal average such that $\langle\hat{\Omega}^{\dagger}_{\rm ad}\hat{\Omega}_{\rm ad}\rangle_{\rm{bare}}<1$, whereas on the effective Hilbert space it holds that,
$\langle\hat{\Omega}_{\rm ad}\hat{\Omega}^{\dagger}_{\rm ad}\rangle_{\rm{eff}}=\langle\hat{\mathbbm{1}}\rangle_{\rm{eff}}=1\;.$ (11)
From Eq. 10 it follows immediately that the minimum is found when the impurity observables match,
@@
for all effective impurity operators $\hat{h}_{i}$ of the effective model (from now on we simplify the notation $\langle\hat{\Omega}_{\rm ad}^{\dagger}\hat{h}_{i}\hat{\Omega}_{\rm ad}\rangle_{\rm{bare}}\rightarrow\langle\hat{h}_{i}\rangle_{\rm bare}$). Thus, in the precise sense of Eq. 9, the optimal low-energy effective model $\hat{H}_{\operatorname{\textrm{eff}}}$ for a given bare model $\hat{H}_{\rm bare}$ is the one that matches the thermodynamic expectation values for all effective impurity operators $\hat{h}_{i}$ in the bare model. We emphasize that the KLD Eq. 9 does not itself ever have to be evaluated. For the GD optimization, it is sufficient to evaluate the gradient of the KLD, which can be expressed in terms of the physical and readily computable observables in Eq. 10. We refer to the process of optimizing an effective model to best represent the low-energy physics of our microscopic bare Hamiltonian by matching these key observables as ‘unsupervised model learning’ (UML).
@@
A particularly attractive feature of UML for quantum impurity systems is the efficiency of the optimization problem. This follows from the fact that the second-order derivative of the KLD corresponds to the second-order derivative of the effective free energy $\mathcal{F}_{\rm eff}(\boldsymbol{\theta})=-1/\beta\log\operatorname{tr}\left[\exp(-\beta\hat{H}_{\rm eff}(\boldsymbol{\theta}))\right]$. Therefore, provided all operators $\{\hat{h}_{i}\}$ mutually commute, the KLD is convex – see Appendix A. Thus we have a single, global minimum, which a gradient descent algorithm is guaranteed to find, making this a trivial optimization problem. We note that the gradient obtained from the KLD, Eq. 10, is equivalent to the gradient obtained in the maximum entropy principle that was recently applied to the problem of Hamiltonian tomography in Ref. [25].
@@
The optimization problem is efficient, given the convexity of Eq. 9. The computational cost of the optimization is therefore mainly controlled by the method used to compute the observables comprising the gradient in Eq. 10. The bottleneck is then the calculation of the observables $\langle\hat{h}_{i}\rangle_{\rm bare}$ in the bare model. However, this needs to be done just once (e.g. with quantum Monte Carlo (QMC) methods [59, 50]). On the other hand, although multiple calculations of $\langle\hat{h}_{i}\rangle_{\rm eff}$ are required during the optimization, these are performed on the simplified effective model and are therefore by construction inexpensive (they can be done with e.g. the numerical renormalization group (NRG) method [51, 52, 53, *stadler2016interleaved]). For the results presented in the following, we have used continuous-time QMC (CT-QMC) and NRG to calculate the observables and the Adagrad or Adam GD methods [60].
II Designing an effective model